OpenStudy (anonymous):

Help please, and maybe an explaination too! What is the solution to the equation? 8 + 4s – 2s = 16 A. s = –12 B. s = –4 C. s = 4 D. s = 12

OpenStudy (shamil98):

Combine the like terms and isolate the variable s.

OpenStudy (shamil98):

4s - 2s + 8 = 16 4s - 2s = 2s 2s + 8 = 16 subtract 8 from both sides 2s = 8 divide by 2 s = 4

OpenStudy (anonymous):

I understand thank you.